# Logarithmic-Depth Quantum Circuits for Hamming Weight Projections

### v1

@@ VI Simulation Results
In this section, we report the results of simulations of our algorithm for a wide range of examples. We benchmarked our algorithm against the true distribution of Hamming weights. For each example, we considered a noiseless, shot noise, and a noisy quantum simulation. Lastly, we also ran the algorithm on actual quantum devices. The results are depicted in Figures 88(t).
@@
In most of the cases, we simulated the width-optimal version of our construction as the depth-optimal version requires a larger number of control qubits and mid-circuit measurements, which are a constraint on most available hardware today. However, for a three-qubit example, we also classically simulate using different $s$ values and the semi-classical encoded IQFT.
@@
All the program code can be found as arXiv ancillary files with the arXiv posting of this paper.
@@
Figure 8: Distribution of counts obtained for an input state $|\psi\rangle=R_{y}\!\left(\frac{\pi}{3}\right)|0\rangle$.
(a) Distribution of counts obtained for an input state $|\psi\rangle=\frac{1}{2}\left(\sqrt{3}|00\rangle+|11\rangle\right)$.
(b) Distribution of counts obtained for a randomly generated two-qubit input state.
@@
(c) Distribution of counts obtained for an input state $|\operatorname{GHZ}\rangle=\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$.
(d) Distribution of counts obtained for an input state $|\operatorname{W}\rangle=\frac{1}{\sqrt{3}}\left(|001\rangle+|010\rangle+|100\rangle\right)$.
(e) Distribution of counts obtained for a randomly generated three-qubit input state.
@@
(f) Distribution of counts obtained for an input state $|\operatorname{GHZ}\rangle=\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$ using $s=1$ and the semi-classical encoded IQFT.
(g) Distribution of counts obtained for an input state $|\operatorname{GHZ}\rangle=\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$ using $s=2$ and the semi-classical encoded IQFT.
(h) Distribution of counts obtained for an input state $|\operatorname{GHZ}\rangle=\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$ using $s=3$ and the semi-classical encoded IQFT.
(i) Distribution of counts obtained for an input state $|\psi\rangle=|\phi^{+}\rangle|\phi^{+}\rangle$.
(j) Distribution of counts obtained for an input state $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|0000\rangle+|1111\rangle\right)$.
@@
(k) Distribution of counts obtained for an input state $|\operatorname{W}_{4}\rangle=\frac{1}{\sqrt{4}}\left(|0001\rangle+|0010\rangle+|0100\rangle+|1000\rangle\right)$.
(l) Distribution of counts obtained for a randomly generated four-qubit input state.
(m) Distribution of counts obtained for an input state $|\psi\rangle=|\operatorname{GHZ}\rangle|\phi^{+}\rangle$.
(n) Distribution of counts obtained for an input state $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|00000\rangle+|11111\rangle\right)$.
@@
(o) Distribution of counts obtained for an input state $|\operatorname{W}_{5}\rangle=\frac{1}{\sqrt{5}}\left(|00001\rangle+|00010\rangle+|00100\rangle+|01000\rangle+|10000\rangle\right)$.
(p) Distribution of counts obtained for a randomly generated five-qubit input state.
(q) Distribution of counts obtained for an input state $|\psi\rangle=|\operatorname{GHZ}\rangle|\operatorname{W}\rangle$.
@@
(r) Distribution of counts obtained for an input state $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|000000\rangle+|111111\rangle\right)$.
(s) Distribution of counts obtained for an input state $|\operatorname{W}_{6}\rangle=\frac{1}{\sqrt{6}}(|000001\rangle+|000010\rangle+|000100\rangle+|001000\rangle+|010000\rangle+|100000\rangle)$.
(t) Distribution of counts obtained for a randomly generated six-qubit input state.
VII Conclusion
@@
Future directions involve proving the optimality of the algorithm and its co-design with various available quantum computational architectures. Let us note here that it can be implemented on a two-dimensional architecture with nearest neighbor interactions, in a manner similar to that outlined in [22]. We will include this two-dimensional implementation in a second version of our paper.
@@ Acknowledgements.
SR thanks Aby Philip and Vishal Singh for helpful and illuminating discussions. SR and MMW acknowledge support from the School of Electrical and Computer Engineering at Cornell University, from the National Science Foundation under Grant No. 2315398, and from AFRL under agreement no. FA8750-23-2-0031. MLL acknowledges support from the DoD SMART Scholarship. This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government.

### v2

@@ VI Simulation Results
In this section, we report the results of simulations of our algorithm for a wide range of examples. We benchmarked our algorithm against the true distribution of Hamming weights. For each example, we considered a noiseless simulation, a simulation with shot noise, and a noisy quantum simulation. Lastly, we also ran the algorithm on actual quantum devices. The results are depicted in Figures 914.
@@
In most of the cases, we simulated the width-optimal version of our construction as the depth-optimal version requires a larger number of control qubits and mid-circuit measurements, which are a constraint on most available hardware today. However, for a three-qubit example, we also classically simulated it using different $s$ values and the semi-classical encoded IQFT.
For the shot-noise simulations, each run used $10,000$ shots. For the noisy simulations, we used the 16-qubit IBM simulator FakeGuadalupeV2. Each run consisted of $10,000$ shots and the median $T_{1}$ and $T_{2}$ values are $7.2\times 10^{-5}$ and $8.8\times 10^{-5}$, respectively. For the quantum device runs, we used the 127-qubit IBM device ibm__\__brisbane. Each run consisted of $10,000$ shots and the median $T_{1}$ and $T_{2}$ values are $2.2\times 10^{-4}$ and $1.3\times 10^{-4}$, respectively.
@@
All the program code is available as arXiv ancillary files along with the arXiv posting of this paper.
@@
Figure 9: Distribution of counts obtained for the input state $|\psi\rangle=R_{y}\!\left(\frac{\pi}{3}\right)|0\rangle$ (top left), the input state $|\psi\rangle=\frac{1}{2}\left(\sqrt{3}|00\rangle+|11\rangle\right)$ (top right), and a randomly generated two-qubit input state (bottom).
@@
Figure 10: Distribution of counts obtained for the input state $|\operatorname{GHZ}\rangle=\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$ (top left), the input state $|\operatorname{W}\rangle=\frac{1}{\sqrt{3}}\left(|001\rangle+|010\rangle+|100\rangle\right)$ (top right), and a randomly generated three-qubit input state (bottom).
@@
Figure 11: Distribution of counts obtained for the input state $|\operatorname{GHZ}\rangle=\frac{1}{\sqrt{2}}\left(|000\rangle+|111\rangle\right)$ using $s=1,2,3$, respectively (top left, top right, bottom), and the semi-classical encoded IQFT.
@@
Figure 12: Distribution of counts obtained for the input state $|\psi\rangle=|\phi^{+}\rangle|\phi^{+}\rangle$ (top left), the input state $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|0000\rangle+|1111\rangle\right)$ (top right), the input state $|\operatorname{W}_{4}\rangle=\frac{1}{\sqrt{4}}\left(|0001\rangle+|0010\rangle+|0100\rangle+|1000\rangle\right)$ (bottom left), and a randomly generated four-qubit input state (bottom right).
@@
Figure 13: Distribution of counts obtained for the input state $|\psi\rangle=|\operatorname{GHZ}\rangle|\phi^{+}\rangle$ (top left), the input state $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|00000\rangle+|11111\rangle\right)$ (top right), the input state $|\operatorname{W}_{5}\rangle=\frac{1}{\sqrt{5}}\left(|00001\rangle+|00010\rangle+|00100\rangle+|01000\rangle+|10000\rangle\right)$ (bottom left), and a randomly generated five-qubit input state (bottom right).
@@
Figure 14: Distribution of counts obtained for the input state $|\psi\rangle=|\operatorname{GHZ}\rangle|\operatorname{W}\rangle$ (top left), the input state $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|000000\rangle+|111111\rangle\right)$ (top right), the input state $|\operatorname{W}_{6}\rangle=\frac{1}{\sqrt{6}}(|000001\rangle+|000010\rangle+|000100\rangle+|001000\rangle+|010000\rangle+|100000\rangle)$ (bottom left), and a randomly generated six-qubit input state (bottom right).
VII Discussion and Conclusion
@@
Future directions involve proving the optimality of the algorithm and its co-design with various available quantum computational architectures. Another future direction is to investigate if the optimal depth scaling is retained when restricted to linear nearest neighbour architectures.
@@ Acknowledgements.
We thank Yingkai Ouyang for helpful comments about using the coherent Hamming weight measurement for quantum error detection and correction. SR thanks Aby Philip and Vishal Singh for helpful and illuminating discussions. SR and MMW acknowledge support from the School of Electrical and Computer Engineering at Cornell University, from the National Science Foundation under Grant No. 2315398, and from AFRL under agreement no. FA8750-23-2-0031. MLL acknowledges support from the DoD SMART Scholarship. This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government.